OpenStudy (anonymous):

m^2-25/2m=5/2 solve by completing the square use radicals as needed use commas to separate answers

OpenStudy (chaise):

m^2 - 25/2m = 5/2 First of all lets rid of the denominators. Multiply both sides by 2m. m^2 - 25 = 5m m^2 -5m - 25 = 0 m^2 - 5m +6.25 -31.25 = 0 (m-(5/2))^2 -31.25 =0 I'm sure you can do the rest. I'm bored.

OpenStudy (chaise):

Unless you need me to do the rest for you and I will help.

OpenStudy (anonymous):

ive been posting most of my answers to see if they were right i couldnt figure this one out

OpenStudy (anonymous):

\[m^2 -25 - 5m = 0\] \[m^2 -25 -5m +25/4 -25/4 = 0\] \[(m - 5/2)^2 = 125/4\] \[m = \frac{5}{2} \pm \sqrt{\frac{125}{4}}\]

OpenStudy (chaise):

Ishaan is correct. I wish I was as good at using the equation editor as him.

OpenStudy (anonymous):

lol

OpenStudy (chaise):

Do you need further explanation or is this fine?